(Homework Assignment 1 (Chapter 2) Problem C2-04 Constants A truck is traveling at a constant speed of 56.3 km/h Suddenly, the driver sees a deer 23.6 m away and slams on the brakes
If the truck hits the deer 2.60 seconds after the driver slams on the brakes, what was the speed of the truck, in km/h, when the deer was struck? Assume motion was at constant acceleration. Answer will be 3 sig figs.
Answer #1

here,

initial speed , u = 56.3 km/h = 15.64 m/s

the distance of deer from the car , s = 23.6 m

let the acceleration be a

time taken , t = 2.6 s

using seccond equation of motion

s = u * t + 0.5 * a * t^2

23.6 = 15.64 * 2.6 + 0.5 * a * 2.6^2

solving for a

a = - 5.048 m/s^2

using first equation of motion

the final speed of car when the deer was struck , v = u + a * t

v = 15.64 - 5.048 * 2.6 m/s

v = 2.51 m/s = 9.04 km/h